# Emergence of Fermi’s Golden Rule in the Probing of a Quantum Many-Body System

### v1

@@ Abstract
The framework of linear response theory is central to our understanding of how to interpret probes of complex quantum systems. For instance, connecting experimental observables to many-body quantities often relies on linear response theory via Fermi’s Golden Rule (FGR) [1, 2]. FGR states that the long-time effect of a perturbation on a many-body system is given by the Fourier transform of the time correlation function of the perturbation operator with the observable. Depending on the context, this can provide crucial microscopic information on the many-body system in equilibrium, such as the spectral function [3, 4], the bulk viscosity [5], and the density susceptibility [6, 7]. The applicability of FGR is wide, ranging from scanning tunneling microscopy [8, 9], optical conductivities [10], to inelastic scattering reactions in nuclear physics [11].
The regime of validity of FGR is a problem of fundamental interest in quantum many-body physics; in strongly correlated systems, determining the validity of FGR a priori is a formidable challenge. FGR generically provides a transition rate from an initial state into a continuum under a coupling perturbation. The three underlying assumptions - a continuum to decay to, a time neither too long nor too short, and a weak coupling — are crucial. Several scenarios where these assumptions are violated have attracted theoretical interest. First, the applicability of FGR to systems with a discrete quasi-continuum was recently studied, e.g. in the context of thermalization of finite-size isolated many-body systems [12]. Secondly, at short times, it is known that FGR does not apply due to unitary time evolution [13, 14]. At long times (i.e. compared to the inverse level spacing of the quasi-continuum), FGR also breaks down - even if first-order perturbation theory might still apply [15]. Thirdly, in some situations, effective couplings might never be weak enough for FGR to apply, even in weakly interacting systems (see e.g. phonons in a weakly interacting 2D superfluid  [16]).
Despite FGR’s pivotal role, there is a scarcity of experimental studies of its applicability. Ultracold gases are attractive platforms to investigate such problems since their Hamiltonians are well known and coupling terms can be accurately engineered [17]. Furthermore, fast timescales associated with out-of-equilibrium many-body dynamics are often directly accessible, see e.g. [18, 19, 20, 21, 22]. Here, we map out the full dynamical response diagram of a strongly correlated quantum many-body system - here a universal spin-1/2 Fermi gas - to a time-dependent momentum-independent coupling, and observe the emergence and breakdown of FGR (Fig. 1).
@@
Conventionally, FGR arises from a linear-response analysis on a system governed by a Hamiltonian $\hat{H}_{\mathrm{S}}$ (which includes both the kinetic and interaction energies) perturbed by a probing term $\hat{V}_{\mathrm{pr}}$. We assume that $\hat{V}_{\mathrm{pr}}$ outcouples particles in spin state $\ket{\downarrow}$ into a noninteracting state |$\ket{\text{o}}$ with driving frequency $\omega_{\mathrm{pr}}$, i.e. $\hat{V}_{\mathrm{pr}}=\frac{\hbar\Omega_{0}}{2}e^{-i\omega_{\mathrm{pr}}t}\sum_{\textbf{k}}\hat{c}_{\text{o},\textbf{k}}^{\dagger}\hat{c}_{\downarrow,\textbf{k}}+\mathrm{h.c.}$, where $\Omega_{0}$ is the bare-transition Rabi frequency characterizing the coupling strength and $\hat{c}_{\sigma,\textbf{k}}$ (resp. $\hat{c}^{\dagger}_{\sigma,\textbf{k}})$ is the annihilation (resp. creation) operator for particles with spin state $\sigma$ and momentum $\hbar\textbf{k}$; $\hat{V}_{\mathrm{pr}}$ does not couple to spin state $\ket{\uparrow}$. The main observable for studying the dynamics induced by $\hat{V}_{\mathrm{pr}}$ is the transition rate $R(\omega,t)\equiv\mathrm{d\alpha}/\mathrm{d}t$ of the fraction of outcoupled particles $\alpha\equiv N_{\text{o}}(\omega,t)/N_{\downarrow}$, where $N_{\downarrow}$ is the initial number of particles in state $\ket{\downarrow}$, $t$ is the coupling pulse duration, and $\omega=\omega_{\mathrm{pr}}-\omega_{0}$ is the detuning to the bare internal-state transition. In a suitable time window, FGR provides a powerful expression for (a time-independent) $R$ [23, 24]:
@@
where $\xi_{k\downarrow}=\hbar^{2}k^{2}/(2m)-\mu_{\downarrow}$, $\textbf{k}\rvert$ k |, $m$ is the atom’s mass, $\mu_{\downarrow}$ is the chemical potential, and $n_{\mathrm{F}}(\epsilon)$ is the Fermi-Dirac distribution at energy $\epsilon$. The most interesting quantity in Eq. (1) is the spectral function $A(\textbf{k},\omega^{\prime})\equiv-2\mathrm{Im}[G_{\mathrm{ret}}(\textbf{k},\omega^{\prime})]$, where $G_{\mathrm{ret}}(\textbf{k},\omega^{\prime})$ is the Fourier transform ($\omega^{\prime}\leftrightarrow t$) of the retarded Green’s function $G_{\mathrm{ret}}(\textbf{k},t)\equiv-i\Theta(t)\braket{\{\hat{c}_{\downarrow,\textbf{k}}(t),\hat{c}_{\downarrow,\textbf{k}}^{\dagger}(0)\}}$, and $\Theta$ is the Heaviside function [25]. It is typically unfeasible to determine a priori the range of $\Omega_{0}$ and $t$ over which Eq. (1) applies in a strongly correlated system.
@@
Figure 3: Dynamics in the linear response regime and emergence of FGR. Time series of the transfer fraction $\alpha(t)$ are shown for various $\Omega_{0}$ and postselected such that $\alpha<0.1$ and $t<12~{}\mathrm{ms}$ [28]. Both in the main panel and the lower inset, the red (resp. gray) dash-dotted line is the universal quadratic $\alpha=\frac{1}{4}\Omega_{0}^{2}t^{2}$ at early times (resp. a fit $\alpha=\frac{Z}{4}\Omega_{0}^{2}t^{2}$, where $Z=0.70(4)$ and the band represents the uncertainty). Upper inset: spectral response $I_{\text{FGR}}(\omega)$ measured in FGR’s regime (see text). Near the peak, we use $\Omega_{0}=2\pi\times 45(3)~{}\mathrm{Hz}$ for $t=5~{}\mathrm{ms}$. Away from the peak, we typically use $\Omega_{0}=2\pi\times 140(5)~{}\mathrm{Hz}$, for $t=1.5~{}\mathrm{ms}$ to improve the signal [28]. The solid line is an interpolation used to determine $\alpha$ via Eq. (2); see the corresponding solid black lines in the main panel and lower inset. As the short-time behavior of $\alpha$ is dictated by the high-$\omega$ part of the rf spectrum, the solid black lines are limited by the range of $\omega$ over which $I_{\mathrm{FGR}}$ was measured (see [28] for a discussion of extrapolations). The dotted blue line in the main panel is determined from the peak spectral response $I_{\text{peak}}$. Lower inset: Early-time dynamics. The ideal-gas data (in practice, a single-component Fermi gas) is shown as red points and a fit gives $\alpha/\Omega_{0}^{2}=0.26(2)\times t^{2.02(5)}$ in excellent agreement with the theoretical limit.
@@
At short times ($E_{\mathrm{F}}t/\hbar\lesssim 3$), the transfer fraction exhibits a $t^{2}$ scaling. At extremely short times, we expect the universal limit $\alpha\cdot(E_{\mathrm{F}}/\hbar\Omega_{0})^{2}\approx(E_{\mathrm{F}}t/\hbar)^{2}/4$, which reflects the normalization of the spectral function [28] (red dot-dashed line). In the lower inset, we show the data taken for the ideal gas (red points), which closely follows that expectation (red dot-dashed line). However, for the shortest times accessed in our work, we find that the interacting-gas measurements are well described by $Z\cdot(E_{\mathrm{F}}t/\hbar)^{2}/4$, with $Z=0.70(4)$ (gray dot-dashed line and band). A well-defined value of $Z$ at intermediate times is indicative of a separation of scales in the spectral response, with a well-defined peak of area $Z$ [31].
@@
The applicability of FGR has important implications for the interpretation of rf spectra measurements [4, 33]. Therefore, we now focus on a more careful examination of the boundary of FGR, in the low-$\Omega_{0}$ regime. We take spectra analogous to the upper inset of Fig. 3 for various $\Omega_{0}$, with a pulse duration $t\approx 2\pi/(5\Omega_{0})$; since we remain close to FGR’s regime, we also normalize the more general spectra to $I\equiv\alpha\cdot(E_{\mathrm{F}}/\hbar\Omega_{0}^{2}t)$. From each of those spectra, we extract the peak frequency shift $\omega_{\mathrm{p}}$, the (full width at half-maximum) width $\Gamma_{\mathrm{fit}}$ and peak spectral response $I_{\mathrm{peak}}$ using a smooth interpolation of the data [28]; we show the results in Fig. 4. Two examples of spectra are shown in the inset of Fig. 4, one within (light red points) and one beyond (dark red points) FGR’s regime for $\omega\approx\omega_{\mathrm{p}}$.
@@
Figure 5: Threshold for Rabi oscillations. (a) Time series $\alpha(t)$ for various $\hbar\Omega_{0}/E_{\mathrm{F}}$. The dash-dotted line and the six-point star indicate the threshold Rabi frequency $\hbar\Omega_{\text{thres}}/E_{\mathrm{F}}$. (b) Oscillation frequency $\Omega_{\mathrm{R}}$ (red points) and decay rate $\Gamma_{\mathrm{R}}$ (blue points) extracted from fits to the time series (see text and [28] for details). The open symbols correspond to $\Omega_{0}\lesssim\Omega_{\text{thres}}$.
@@
Figure 6: Dynamical response diagram of the universal spin-$1/2$ Fermi gas versus the coupling control parameters $\Omega_{0}$ and $t$. The dotted lines denote the (non-sharp) boundaries between regimes; $\lambda$ and $\Gamma_{0}$ are FGR’s transition rate and spectral width. The stars represent the rescaled data from Fig. 2, illustrating the condition where $\alpha=10\%$. Diamond and square markers, which correspond to various measurements from Figs. 3-4, with matching colors, highlight the boundaries between different regimes. The dash-dotted line and six-point star indicate the threshold frequency for Rabi oscillations as in Fig. 5, and the band represents the uncertainty.
@@
For large enough drives, the non-perturbative dynamics will ultimately consist of Rabi oscillations, which can exhibit many-body effects [34, 35, 22]. We now examine the emergence of those oscillations. In Fig. 5a we show a map of the dynamics of $\alpha(t)$ as a function of $t$ and $\Omega_{0}$. At low $\Omega_{0}$, $\alpha(t)$ is mostly monotonically increasing. For $\Omega_{0}$ larger than a threshold $\Omega_{\text{thres}}$, Rabi oscillations appear, with an oscillation frequency $\Omega_{\text{R}}\sim\Omega_{0}$; we determine that $\hbar\Omega_{\text{thres}}/E_{\mathrm{F}}=0.17(2)$ (see time series in [28]). Calculating this threshold theoretically is a complex out-of-equilibrium many-body problem. However, we remark that $\Omega_{\text{thres}}/\Gamma_{0}=0.7(2)$, which suggests, somewhat analogously to the atom-photon case [32], that this transition occurs in a many-body system when the Rabi frequency is comparable to the continuum width. Recent measurements on the highly spin-imbalanced Fermi gas [22] (i.e. $N_{\downarrow}\ll N_{\uparrow}$) also support this idea; indeed, from the data in [22], we extract $\Omega_{\text{thres}}/\Gamma_{0}=0.8(2)$.
@@
For $\Omega_{0}>\Omega_{\text{thres}}$, we extract the oscillation frequency $\Omega_{\mathrm{R}}$ and the decay rate $\Gamma_{\mathrm{R}}$ from a fit, see Fig. 5b. The renormalized oscillation frequency $(\Omega_{\mathrm{R}}/\Omega_{0})^{2}$ shows a plateau at low $\Omega_{0}$, down to $\Omega_{0}\approx\Omega_{\text{thres}}$. Its plateau value $(\Omega_{\mathrm{R}}/\Omega_{0})^{2}=0.73(5)$ is close to an analogous measurement in the highly-imbalanced Fermi gas, $0.63(3)$ [22, 36]. For $\hbar\Omega_{0}\gtrsim E_{\mathrm{F}}$, $\Omega_{\mathrm{R}}/\Omega_{0}$ rises rapidly. At high $\Omega_{0}$, $\Omega_{\mathrm{R}}$ approaches $\Omega_{0}$; in that regime, $\Omega_{0}/\Gamma_{0}\gtrsim 15$, so that the many-body system can be approximated as a two-level system [37, 32]. At low $\Omega_{0}$, $\Gamma_{\mathrm{R}}$ decreases with $\Omega_{0}$, without reaching a nonzero limit. At $\Omega_{0}\approx 4E_{\mathrm{F}}/\hbar$, $\Gamma_{\mathrm{R}}$ approaches a maximum, $\hbar\Gamma_{\mathrm{R}}\approx 4E_{\mathrm{F}}$. The analogous measurements in the highly-imbalanced system were connected to many-body properties of the dressed (Fermi polaron) impurity, e.g. the quasiparticle residue and its lifetime [22]. Intriguingly, even though such simple interpretations are elusive in this spin-balanced system (as there is no small parameter, and no well-defined quasiparticles in the normal phase), the results are qualitatively similar [28, 38].
@@
We thank S. Girvin and L. Glazman for fruitful discussions, and R. Fletcher, G. Martirosyan, M. Gemaljevic, and E. Papoutsis for comments on the manuscript. We thank M. Zwierlein and B. Mukherjee for sharing their data. This work was supported by the AFOSR (Grant No. FA9550-23-1-0605), NSF (Grant No. PHY-1945324), the David and Lucile Packard Foundation, and the Alfred P. Sloan Foundation. G.L.S. acknowledges support from the NSF Graduate Research Fellowship Program. A.S. acknowledges support from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator.
Supplemental Material

### v2

@@ Abstract
The framework of linear response theory is central to our understanding of how to interpret probes of complex quantum systems. For instance, connecting experimental observables to many-body quantities often relies on linear response theory via Fermi’s Golden Rule (FGR) [1, 2]. FGR states that the long-time effect of a perturbation on a many-body system is given by the Fourier transform of the time correlation function of the perturbation operator with the observable. Depending on the context, this can provide crucial microscopic information on the many-body system in equilibrium, such as spectral functions, dynamic structure factors or local density of states (see e.g. [3, 4, 5]). The applicability of FGR is wide, ranging from describing scanning tunneling microscopy [6, 7], optical conductivity measurements [8], to inelastic scattering reactions in nuclear physics [9].
The regime of validity of FGR is a problem of fundamental interest in quantum many-body physics; in strongly correlated systems, determining the validity of FGR a priori is a formidable challenge. FGR generically provides a transition rate from an initial state into a continuum under a coupling perturbation. The three underlying assumptions - a continuum to decay to, a time neither too long nor too short, and a weak coupling — are crucial. Several scenarios where these assumptions are violated have attracted theoretical interest. First, the applicability of FGR to systems with a discrete quasi-continuum was recently studied, e.g. in the context of thermalization of finite-size isolated many-body systems (see e.g. [10, 11, 12]). Secondly, at short times, it is known that FGR does not apply due to unitary time evolution [13, 14]. At long times (i.e. compared to the inverse level spacing of the quasi-continuum), FGR also breaks down - even if first-order perturbation theory might still apply [15]. Thirdly, in some situations, effective couplings might never be weak enough for FGR to apply, even in weakly interacting systems (see e.g. phonons in a weakly interacting 2D superfluid  [16]).
Despite FGR’s pivotal role, there is a scarcity of experimental studies of its applicability. Ultracold gases are attractive platforms to investigate such problems since their Hamiltonians are well known and coupling terms can be accurately engineered [17]. Furthermore, fast timescales associated with out-of-equilibrium many-body dynamics are often directly accessible, see e.g. [18, 19, 20, 21, 22]. Here, we map out the full dynamical response diagram of a strongly correlated quantum many-body system - a universal spin-1/2 Fermi gas - to a time-dependent momentum-independent coupling, and observe the emergence and breakdown of FGR (Fig. 1).
@@
Traditionally, FGR arises from a linear response analysis on a system governed by a Hamiltonian $\hat{H}_{\mathrm{S}}$ perturbed by a probing term $\hat{V}_{\mathrm{pr}}$. We assume that $\hat{V}_{\mathrm{pr}}$ is turned on at $t=0$ and outcouples particles in spin state $\ket{\downarrow}$ into a noninteracting state |$\ket{\text{o}}$ with driving frequency $\omega_{\mathrm{pr}}$ as $\hat{V}_{\mathrm{pr}}=\hbar\Omega_{0}\cos\left(\omega_{\mathrm{pr}}t\right)\sum_{\textbf{k}}\hat{c}_{\text{o},\textbf{k}}^{\dagger}\hat{c}_{\downarrow,\textbf{k}}+\mathrm{h.c.}$, where $\Omega_{0}$ is the bare-transition Rabi frequency characterizing the coupling strength and $\hat{c}_{\sigma,\textbf{k}}$ (resp. $\hat{c}^{\dagger}_{\sigma,\textbf{k}})$ is the annihilation (resp. creation) operator for particles with spin state $\sigma$ and momentum $\hbar\textbf{k}$; note that $\hat{V}_{\mathrm{pr}}$ does not couple to spin state $\ket{\uparrow}$. The main observable for studying the dynamics induced by $\hat{V}_{\mathrm{pr}}$ is the transition rate $R(\omega,t)\equiv\mathrm{d\alpha}/\mathrm{d}t$ of the fraction of outcoupled particles $\alpha\equiv N_{\text{o}}(\omega,t)/N_{\downarrow}$, where $N_{\downarrow}$ is the initial number of particles in state $\ket{\downarrow}$, $t$ is the coupling pulse duration, and $\omega=\omega_{\mathrm{pr}}-\omega_{0}$ is the detuning to the bare internal-state transition. In a suitable time window, FGR provides a powerful expression for (a time-independent) $R$ [23, 24]:
@@
where $\xi_{k\downarrow}=\hbar^{2}k^{2}/(2m)-\mu_{\downarrow}$, $\textbf{k}\rvert$ k |, $m$ is the atom’s mass, $\mu_{\downarrow}$ is the chemical potential, and $n_{\mathrm{F}}(\epsilon)$ is the Fermi-Dirac distribution at energy $\epsilon$. The most interesting quantity in Eq. (1) is the spectral function $A(\textbf{k},\omega^{\prime})$, which encodes information on the excitations and the density of states of a many-body system [4, 25].
@@
Figure 3: Dynamics in the linear response regime and emergence of FGR. Time series of the transfer fraction $\alpha(t)$ are shown for various $\Omega_{0}$ and postselected such that $\alpha<0.1$ and $t<12~{}\mathrm{ms}$ [25]. Both in the main panel and the lower inset, the red (resp. gray) dash-dotted line is the universal quadratic $\alpha=\frac{1}{4}\Omega_{0}^{2}t^{2}$ at early times (resp. a fit $\alpha=\frac{Z}{4}\Omega_{0}^{2}t^{2}$, where $Z=0.70(4)$ and the band represents the uncertainty). Upper inset: spectral response $I_{\text{FGR}}(\omega)$ measured in FGR’s regime (see text). Near the peak, we use $\Omega_{0}=2\pi\times 45(3)~{}\mathrm{Hz}$ for $t=5~{}\mathrm{ms}$. Away from the peak, we typically use $\Omega_{0}=2\pi\times 140(5)~{}\mathrm{Hz}$, for $t=1.5~{}\mathrm{ms}$ to improve the signal [25]. The solid line is an interpolation used to determine $\alpha$ via Eq. (2); see the corresponding solid black lines in the main panel and lower inset. As the short-time behavior of $\alpha$ is dictated by the high-$\omega$ part of the rf spectrum, the solid black lines are limited by the range of $\omega$ over which $I_{\mathrm{FGR}}$ was measured (see [25] for a discussion of extrapolations). The dotted blue line in the main panel is determined from the peak spectral response $I_{\text{peak}}$. Lower inset: Early-time dynamics. The ideal-gas data (in practice, a single-component Fermi gas) is shown as red points and a fit gives $\alpha/\Omega_{0}^{2}=0.26(2)\times t^{2.02(5)}$, in excellent agreement with the theoretical limit.
@@
At short times ($E_{\mathrm{F}}t/\hbar\lesssim 3$), the transfer fraction exhibits a $t^{2}$ scaling. At extremely short times, we expect the universal limit $\alpha\cdot(E_{\mathrm{F}}/\hbar\Omega_{0})^{2}\approx(E_{\mathrm{F}}t/\hbar)^{2}/4$, which reflects the normalization of the spectral function [25] (red dot-dashed line). In the lower inset, we show the data taken for the ideal gas (red points), which closely follows that expectation (red dot-dashed line). However, for the shortest times accessed in our work, we find that the interacting-gas measurements are instead well described by $Z\cdot(E_{\mathrm{F}}t/\hbar)^{2}/4$, with $Z=0.70(4)$ (gray dot-dashed line and band). A well-defined value of $Z$ at intermediate times is indicative of a separation of scales in the spectral response, with a well-defined peak of area $Z$ [30].
@@
The applicability of FGR has important implications for the interpretation of rf spectra measurements [32, 33] (see also [34] for the measurement of the same spectrum using a microwave transition). Therefore, we now focus on a more careful examination of the boundary of FGR, in the low-$\Omega_{0}$ regime. We take spectra analogous to the upper inset of Fig. 3 for various $\Omega_{0}$, with a pulse duration $t\approx 2\pi/(5\Omega_{0})$; since we remain close to FGR’s regime, we also normalize the more general spectra to $I\equiv\alpha\cdot(E_{\mathrm{F}}/\hbar\Omega_{0}^{2}t)$. From each of those spectra, we extract the peak frequency shift $\omega_{\mathrm{p}}$, the (full width at half-maximum) width $\Gamma_{\mathrm{fit}}$ and peak spectral response $I_{\mathrm{peak}}$ using a smooth interpolation of the data [25]; we show the results in Fig. 4. Two examples of spectra are shown in the inset of Fig. 4, one within (light red points) and one beyond (dark red points) FGR’s regime for $\omega\approx\omega_{\mathrm{p}}$.
@@
Figure 5: Threshold for Rabi oscillations. (a) Time series $\alpha(t)$ for various $\hbar\Omega_{0}/E_{\mathrm{F}}$. The dash-dotted line and the six-point star indicate the threshold Rabi frequency $\hbar\Omega_{\text{thres}}/E_{\mathrm{F}}$. (b) Oscillation frequency $\Omega_{\mathrm{R}}$ (light red points) and decay rate $\Gamma_{\mathrm{R}}$ (blue points) extracted from fits to the time series (see text and [25] for details). The open symbols correspond to $\Omega_{0}\lesssim\Omega_{\text{thres}}$.
@@
Figure 6: Dynamical response diagram of the universal spin-$1/2$ Fermi gas versus the coupling control parameters $\Omega_{0}$ and $t$. The dotted lines mark the (non-sharp) boundaries between regimes; $\lambda$ and $\Gamma_{0}$ are FGR’s transition rate and spectral width. The stars represent the rescaled data from Fig. 2, illustrating the condition where $\alpha=10\%$. Diamond and square markers, which correspond to various measurements from Figs. 3-4, with matching colors, highlight the boundaries between different regimes. The dash-dotted line and six-point star indicate the threshold frequency for Rabi oscillations as in Fig. 5, and the band represents the uncertainty.
@@
For large enough drives, the non-perturbative dynamics will ultimately consist of Rabi oscillations, which can exhibit many-body effects [35, 36, 22]. We now examine the emergence of those oscillations. In Fig. 5a we show a map of the dynamics of $\alpha(t)$ as a function of $t$ and $\Omega_{0}$. At low $\Omega_{0}$, $\alpha(t)$ is mostly monotonically increasing. For $\Omega_{0}$ larger than a threshold $\Omega_{\text{thres}}$, Rabi oscillations appear, with an oscillation frequency $\Omega_{\text{R}}\sim\Omega_{0}$; we determine that $\hbar\Omega_{\text{thres}}/E_{\mathrm{F}}=0.17(2)$ (see the time series data in [25]). Calculating this threshold theoretically is a complex out-of-equilibrium many-body problem. However, we remark that $\Omega_{\text{thres}}/\Gamma_{0}=0.7(2)$, which suggests, somewhat analogously to the atom-photon case [31], that this transition occurs in a many-body system when the Rabi frequency is comparable to the continuum width. Recent measurements on the highly spin-imbalanced Fermi gas [22] (i.e. $N_{\downarrow}\ll N_{\uparrow}$) also support this idea; indeed, from the data in [22], we determine that $\Omega_{\text{thres}}/\Gamma_{0}=0.8(2)$.
@@
For $\Omega_{0}>\Omega_{\text{thres}}$, we extract the oscillation frequency $\Omega_{\mathrm{R}}$ and the decay rate $\Gamma_{\mathrm{R}}$ from a fit, see Fig. 5b. The renormalized oscillation frequency $(\Omega_{\mathrm{R}}/\Omega_{0})^{2}$ shows a plateau at low $\Omega_{0}$, down to $\Omega_{0}\approx\Omega_{\text{thres}}$. Its plateau value $(\Omega_{\mathrm{R}}/\Omega_{0})^{2}=0.73(5)$ is close to an analogous measurement in the highly-imbalanced Fermi gas, $0.63(3)$ [22, 37]. For $\hbar\Omega_{0}\gtrsim E_{\mathrm{F}}$, $\Omega_{\mathrm{R}}/\Omega_{0}$ rises rapidly. At high $\Omega_{0}$, $\Omega_{\mathrm{R}}$ approaches $\Omega_{0}$; in that regime, $\Omega_{0}/\Gamma_{0}\gtrsim 15$, so that the many-body system can be approximated as a two-level system [38, 31]. At low $\Omega_{0}$, $\Gamma_{\mathrm{R}}$ decreases with $\Omega_{0}$, without reaching a nonzero limit. At $\Omega_{0}\approx 4E_{\mathrm{F}}/\hbar$, $\Gamma_{\mathrm{R}}$ approaches a maximum, $\hbar\Gamma_{\mathrm{R}}\approx 0.8E_{\mathrm{F}}$. The analogous measurements in the highly-imbalanced system were connected to many-body properties of the dressed (Fermi polaron) impurity, e.g. the quasiparticle residue and its lifetime [22]. Intriguingly, even though such simple interpretations are elusive in this spin-balanced system (as there is no small parameter, and no well-defined quasiparticles in the normal phase), the results are qualitatively similar [25, 39].
@@
We thank S. Girvin and L. Glazman for fruitful discussions, and R. Fletcher, G. Martirosyan, M. Gemaljevic, and E. Papoutsis for comments on the manuscript. We thank M. Zwierlein and B. Mukherjee for sharing their data. This work was supported by the NSF (Grant No. PHY-1945324), AFOSR (Grant No. FA9550-23-1-0605), the David and Lucile Packard Foundation, and the Alfred P. Sloan Foundation. G.L.S. acknowledges support from the NSF Graduate Research Fellowship Program. A.S. acknowledges support from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator.
Supplementary information